Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day07


Crew 97 Log Book for January 7, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status:
Good.

Time Departed/Returned from EVA:
EVA #7: 09:34 / 10:53. 2 Crewmembers
EVA #8: 14:16 / 14:58. 2 Crewmembers
EVA #9: 15:34 / 16:42. 1 Crewmember
EVA #10: 17:05 / 17:50. 2 Crewmembers

Brief Narrative of Field Mission Results:
Busy day of EVAs to collect satellite data, analyze water gullies, and successfully operate and analyze the rover on EVA.

EVA Data/Interpretations:

EVA #7: Samples successfully taken for Satellite Comparison study.

EVA #8: Attempted to study Water Gully slush flow, but incline was too steep to climb.

EVA #9: Tested "Sandstorm" Rover. Rover was controlled inside the Hab and operated in the outside area near the Hab along with EVA crew-member.

EVA #10: Crew-members followed the tracks which the rover traversed in order to analyze the capabilities of the rover.

Engineering/Hab Maintenance:
- Toilet flush line now operational after installation of a new bladder.
- Working on repairing hot water heater.
- Painted black the tank cover area of the outside potable water tank.
- Hab power outages in the late afternoon. Power successfully restored.
- Pack 4 repaired by repairing hoses with new caulk.

Report Transmission Schedule:
CMDR Check-in, ENGR, CMDR Report, Photos, 4 EVA Reports, 3 Science Reports

Plans for Tomorrow:
- 07:30 - Breakfast (2 Crewmembers)
- 8:00 - Crew exercise
- 8:30 - Psychophysiological Experiment
- Engineering checks/repairs
- EVA #11: Water Gully
- 11:30 - Lunch
- Psychophysiological Experiment
- Engineering
- EVA #12: Rock Varnish
- Crew experiments
- Engineering checks. Repair pack 3
- Possible rover EVA
- 17:30 - Dinner
- Psychophysiological Experiment
- Reports by 20:00
- Food Study Questionnaire
- Work on experiments
- Showers: 2 Crewmembers

Inventory:
n/a

Support Requested:
n/a

Miscellaneous:
n/a


Commander's Report

Judah Epstein Reporting

Despite power outages, today was a very busy and productive days on Mars with multiple EVAs, scientific experiments, and engineering repairs.

After breakfast and exercise, the morning began with the first full day of psychophysiological analysis followed by an EVA to collect satellite data comparison. The early afternoon EVA consisted of studying a water gully.

Two additional late afternoon EVAs were added to the schedule because of the successful in-Hab operation of the rover! Therefore an EVA was created to test the rover in outside Martian conditions. A follow-up EVA took place without the rover, but for Marsonauts to follow the tracks which the rover took. Upon return to the Hab, the crew was delayed in the airlock due to a power outage, and thus remained in the airlock for a short time while the Crew Engineer reinstated power to the Hab and the repressurization sequence.

Engineering was busy throughout the day to return the water heater to operation. The system was extensively analyzed and it was determined that the water heater was receiving power, but the gas does not seem to be functioning to the system. Additionally the outside potable water tank cover area was painted black in order to retain additional heat to reduce potential for freezing of the water. Furthermore, in the late afternoon the Hab encountered several power failures. Fortunately all that was required was to reset the outside breaker at the generator. Current drawing systems in the Hab were reduced in order to maintain power to essential equipment.

Tomorrow will consist of extended psycho-physiological analysis, extended water filtration testing including bacteria results, as well as EVAs to study Rock Varnish, Water Gullies, and possibly another Martian Rover EVA.

Science Report 1

Judah Epstein Reporting

Despite power outages, today was a very busy and productive days on Mars with multiple EVAs, scientific experiments, and engineering repairs.

After breakfast and exercise, the morning began with the first full day of psychophysiological analysis followed by an EVA to collect satellite data comparison. The early afternoon EVA consisted of studying a water gully.

Two additional late afternoon EVAs were added to the schedule because of the successful in-Hab operation of the rover! Therefore an EVA was created to test the rover in outside Martian conditions. A follow-up EVA took place without the rover, but for Marsonauts to follow the tracks which the rover took. Upon return to the Hab, the crew was delayed in the airlock due to a power outage, and thus remained in the airlock for a short time while the Crew Engineer reinstated power to the Hab and the repressurization sequence.

Engineering was busy throughout the day to return the water heater to operation. The system was extensively analyzed and it was determined that the water heater was receiving power, but the gas does not seem to be functioning to the system. Additionally the outside potable water tank cover area was painted black in order to retain additional heat to reduce potential for freezing of the water. Furthermore, in the late afternoon the Hab encountered several power failures. Fortunately all that was required was to reset the outside breaker at the generator. Current drawing systems in the Hab were reduced in order to maintain power to essential equipment.

Tomorrow will consist of extended psycho-physiological analysis, extended water filtration testing including bacteria results, as well as EVAs to study Rock Varnish, Water Gullies, and possibly another Martian Rover EVA.

Science Report 2

Tonya Thompson Reporting

I performed 3 physiological reading sessions on each crewmember today.

Each session lasted 3 minutes.

The crewmember entered my stateroom, sat in a chair, placed finger sensors on their 3 middle fingers, and watched a white image projected onto a wall for 3 minutes.

During that time, I recorded their heart rate variability (hrv) and skin conductance level (scl).

The data for each session was saved to the hard drive of a desktop pc in my stateroom.

I also had each crewmember fill out a Cohen’s Perceived Stress Scale survey at the end of the evening after dinner.

Science Report 3

Lucinda Land Reporting

We had a successful first day starting up Sandstorm, now lovingly called "Sandy" by a Crew 97 member. All systems were nominal. Our XO drove Sandstorm around the first floor of the Hab otherwise known as the lab. Driving the rover consists of using a joystick and watching a monitor to tell where it is heading. As a test of this and his abilities to drive the rover, we moved Sandstorm to the second floor of the Hab and our XO stayed on the first floor where he was remotely driving it. He drove it around the living quarters and into the Commander's room to experience its maneuverability.

Three main components are needed to drive the rover, a laptop or monitor with keyboard, a router, and a joystick (keyboard controls are possible as well). We needed to test the range of the router with the rover. The router signal range has a limit and is how the rover receives commands from the joystick. We had no problems with the rover receiving a signal anywhere inside the Hab. The XO suited up and decided to do an EVA with the rover to test the router signal strength reaching outside the Hab.

First the XO took Sandstorm inside the airlock and we tested the signal strength in there by driving it around. It was good. Next the XO took the rover outside the Hab and put Sandstorm on the level ground with him. The rover was still receiving signals from the router inside the Hab. The XO and the rover were out on an EVA for over an hour as he followed Sandstorm west of the Hab, Sandstorm went a total of 193.12 meters. It is not conclusive yet, but it is thought that that rover's batteries died before the router's signal strength lessened. Sandstorm's batteries last on average about one hour.

It was a good day for Sandstorm and her tests. Next, we will test her GPS logging, mapping, snapshot, and Gigapan imagery capabilities. More tests to come. These tests will take place on further EVA's outside the Hab.

Engineering Report

Nathan Wong Reporting


Generator/Electricity:

Kitty(Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): Yes

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
Power continually went out in the hab, we originally thought that it was because of the hot water heater switch, but power kept going off even if the switch was off. Power is restored and now working well.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): .75

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.57

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 4 Gas cans full

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 3

Honey Oil Quantity (# of Quarts in storage): 2

ATV Oil Quantity (# of Quarts in storage): 2

Car Oil Quantity (# of Quarts in storage): none


Water Status:

Outside Potable Water Tank Level (inches from bottom): Half

Trailer Potable Water Tank Level (inches from bottom): Full

Hab Potable Water Tank Level (inches from bottom): Full

Potable Water Meter Reading: 40618.3

In to GreenHab Meter Reading:

Notes/Comments:


GreenHab:

Condition of plants in Tank 1: None

Condition of plants in Tank 2: 3 plants

Condition of plants in Tank 3: 7 plants

Greenhab Notes/Comments: Plants look good.


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments: Observatory not used


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: good
Odometer at end of day: 123085
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments: ATV 1 does not idle, and needs to have the battery charged every morning in order to function properly.

ATV 2 (Spirit):
Used: Yes
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments: Working well.

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments: ATV 3 is possible to start with a pull cord, but this is very hard to do while on EVA.

ATV 4 (Viking II):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments: Working well.


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): off
Actual temperature upstairs (Farenheit): 72

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 57


Computers/Networking Infrastructure:

Notes/comments on internet/computers: Everything is working well


General Engineering Notes/Comments:

-Continued to working on the flush line, a new bladder was installed and it helped the pressure.
-Repairs were made on pack 4. The hoses kept falling out, new caulk was applied
-Pack 3 is still not operational
-The water tank was painted black to absorb more solar energy.
-Started to look at the hot water heater, could not fix.

EVA 7 Report

Jim Crowell Reporting

Time: 09:34 / 10:53

Crew Members: Amanda Damptz and Jim Crowell

Site Location: South of the Hab, along the ridge

Transit Mode: Walking

Objectives: To reach a point along the ridge to the South of the Hab and collect panorama images

Results: Point desired reached. Five samples taken; three on the way to final location, one at final location, and one on the return to the Hab. Also - slush flow location found for crewmember experiment.

Lessons Learned: Sunglasses are not necessary in all cases, and they can decrease visibility greater, making them more of a hindrance than a benefit.


EVA 8 Report

Jim Crowell Reporting

Time: 14:16 / 14:58

Crew Members: Amanda Damptz, Judah Epstein, Jim Crowell

Site Location: South of the Hab, back to where slush flow was found in EVA 7

Transit Mode: ATVs

Objectives: To measure the height of the slush flow

Results: Ascent to the top of the flow from one location was too steep, making it impossible to reach. It was decided to abort the EVA and return on foot at a later time.

Lessons Learned: Quick access to an area is not always the best route to take.


EVA 9 Report

Jim Crowell Reporting

Time: 15:34 / 16:42

Crew Members: Jim Crowell

Site Location: West of the Hab, along the road

Transit Mode: Walking

Objectives: To test the Sensetta J Class rover, known as "Sandstorm," with the controller (Lucinda Land) inside the Hab

Results: The rover was able to travel a distance of 193.12 metres away from the Hab, across dirt, mud, slush, snow, and ice. The rover lost its ability to communicate with the controller behind hills. Near the end of the EVA, the batteries in the rover lost power, causing the rover to stop functioning.

Lessons Learned: The rover can travel farther away from the controller than originally anticipated, and it can operate with the controller inside the Hab.

EVA 10 Report

Jim Crowell Reporting

Time: 17:05 / 17:50

Crew Members: Lucinda Land, Jim Crowell

Site Location: West of the Hab, same as EVA 9

Transit Mode: Walking

Objectives: To analyze and photograph the track the Sensetta J Class rover "Sandstorm" took during EVA 9

Results: Tracks of "Sandstorm" were easily visible and photographs taken successfully.

Lessons Learned: (the controller) was able to see the terrain the tracks crossed during EVA 9, allowing her to make better assessments of what the rover can handle during future interactions.